# Spin-resolved nonlocal transport in proximitized Rashba nanowires

### v1

@@ IV Conclusions
We analyzed in detail spin and transport properties of a proximitized Rashba nanowire in a three-terminal setup with grounded superconductor and with tunnel-coupled side normal or spin polarized leads. We have considered several configurations of the spin polarization of the leads, revealing a distinct correspondence between the sign of the nonlocal conductance and the sign of the
local quasiparticle spin density. In particular, in the setups featuring two leads with the same or opposite spin polarizations, we observed a dominance of either the ECT or CAR processes, if the spin of the probed quasiparticle state matches the spin of the injected charge carrier in the lead. Alternatively, employing one normal and one spin-polarized lead facilitated the precise mapping between the sign of the nonlocal conductance and the spin polarization of the probed quasiparticle state. Furthermore, nonlocal conductances in setups with two normal leads provide information solely pertaining to the quasiparticle charge, rather than spin. We showed that such a behavior can be used to detect the TPT which involves band inversion and a related sign inversion of spin and charge of the lowest energy states. Moreover, we showed that a Rashba nanowire with tunnel-coupled side leads can be a versatile platform for tuning between CAR and ECT processes. This functionality holds promise for applications such as Cooper pair splitters and minimal Kitaev chain systems of quantum dots hosted in proximitized Rashba nanowires. Importantly, our findings highlight the importance of coupling leads to regions away from the nanowire ends, which gives additional insight into the charge and spin density characteristics of quasiparticles. We expect that our results will be particularly useful to experimentalists working on hybrid superconductor-semiconductor systems.
@@ Appendix B Shorter Wire Limit
In this Appendix, we consider the setup with left lead being spin up/down polarized and the right lead being normal. The corresponding nonlocal conductances $G_{L(\pm)R(N)}(\Delta_{Z},E)$ for the nanowire of the length $L=3$ $\mu$m ($N=100$, $a=30$ nm, $t\approx 3$ meV, $t_{\Gamma}=0.15t\approx 0.45$ meV, $x_{L}/a=45$, $x_{R}/a=55$) and $L=1$ $\mu$m ($N=100$, $a=10$ nm, $t\approx 27.21$ meV, $t_{\Gamma}=0.15t\approx 4.1$ meV, $x_{L}/a=45$, $x_{R}/a=55$) are presented on Fig. 11(a,b) and Fig. 12(a,b), respectively. Here, we choose the distances between the gates as $d=x_{R}-x_{L}=300$ nm and $d=x_{R}-x_{L}=100$ nm. The corresponding energy spectrum calculated as a function of Zeeman energy $\Delta_{Z}$ for the proximitized nanowire is shown in Fig. 11 (c,d) and
@@ Appendix C Lattice spacing, numerical check
As we mentioned in the main text, for the purpose of numerical efficiency, we set the effective tight binding lattice spacing to $a=100$ nm in order to study the long wire limit ($L=10$ $\mu$m with $N=100$). Here, we show that, for $10$ times smaller lattice spacing $a=10$ nm ($L=10$ $\mu$m with $N=1000$, $t\approx 27.21$ meV, $t_{\Gamma}=0.2t\approx 5.45$ meV), the key results are almost identical. As an example, we consider the setup with left lead being spin-up polarized and right lead being unpolarized and compare directly the corresponding nonlocal conductance map $G_{L(+)R_{(}N)}(\Delta_{Z},E)$ for the parameters $a=100$ nm ($N=100$) and $a=10$ nm ($N=1000$), see Fig. 13.
Figure 13: The comparison of the nonlocal conductance $G_{L(+)R(N)}(\Delta_{Z},E)$ obtained for different lattice parameters for the same nanowire length $L=10$ $\mu$m. Results for $a=100$ nm and $N=100$ [as in the main text, see Fig. 7(a)] and for $a=10$ nm and $N=1000$ are presented on panel (a) and (a’), respectively. In the regime of interest, $|E|<\Delta_{sc}$, the obtained results are almost identical.

### v2

@@ IV Conclusions
We analyzed in detail spin and transport properties of a proximitized Rashba nanowire in a three-terminal setup with grounded superconductor and with tunnel-coupled side normal or spin polarized leads. We have considered several configurations of the spin polarization of the leads, revealing a distinct correspondence between the sign of the nonlocal conductance and the sign of the local quasiparticle spin density. In particular, in the setups featuring two leads with the same or opposite spin polarizations, we observed a dominance of either the ECT or CAR processes, if the spin of the probed quasiparticle state matches the spin of the injected charge carrier in the lead. Alternatively, employing one normal and one spin-polarized lead facilitated the precise mapping between the sign of the nonlocal conductance and the spin polarization of the probed quasiparticle state. Furthermore, nonlocal conductances in setups with two normal leads provide information solely pertaining to the quasiparticle charge, rather than spin. We showed that such a behavior can be used to detect the TPT which involves band inversion and a related sign inversion of spin and charge of the lowest energy states. Moreover, we showed that a Rashba nanowire with tunnel-coupled side leads can be a versatile platform for tuning between CAR and ECT processes. This functionality holds promise for applications such as Cooper pair splitters and minimal Kitaev chain systems of quantum dots hosted in proximitized Rashba nanowires. Importantly, our findings highlight the importance of coupling leads to regions away from the nanowire ends, which gives additional insight into the charge and spin density characteristics of quasiparticles. We expect that our results will be particularly useful to experimentalists working on hybrid superconductor-semiconductor systems.
@@ Appendix B Shorter Wire Limit
In this Appendix, we consider the setup with left lead being spin up/down polarized and the right lead being normal. The corresponding nonlocal conductances $G_{L(\pm)R(N)}(\Delta_{Z},E)$ for the nanowire of the length $L=3$ $\mu$m ($N=100$, $a=30$ nm, $t\approx 3$ meV, $t_{\Gamma}=0.15t\approx 0.45$ meV, $x_{L}/a=45$, $x_{R}/a=55$) and $L=1$ $\mu$m ($N=100$, $a=10$ nm, $t\approx 27.21$ meV, $t_{\Gamma}=0.15t\approx 4.1$ meV, $x_{L}/a=45$, $x_{R}/a=55$) are presented in Fig. 12(a,b) and Fig. 13(a,b), respectively. Here, we choose the distances between the gates as $d=x_{R}-x_{L}=300$ nm and $d=x_{R}-x_{L}=100$ nm. The corresponding energy spectrum calculated as a function of Zeeman energy $\Delta_{Z}$ for the proximitized nanowire is shown in Fig. 12 (c,d) and
@@
Figure 14: The comparison of the nonlocal conductance $G_{L(+)R(N)}(\Delta_{Z},E)$ obtained for different lattice parameters for the same nanowire length $L=10$ $\mu$m. Results for $a=100$ nm and $N=100$ [as in the main text, see Fig. 7(a)] and for $a=10$ nm and $N=1000$ are presented on panel (a) and (a’), respectively. In the regime of interest, $|E|<\Delta_{sc}$, the obtained results are almost identical. Panels (b) and (b’) illustrate nonlocal conductances for systems with disorder, which is depicted correspondingly in Fig. 3(d) and Fig. 15(c).
@@ Appendix C Lattice spacing, numerical check
As we mentioned in the main text, for the purpose of numerical efficiency, we set the effective tight binding lattice spacing to $a=100$ nm in order to study the long wire limit ($L=10$ $\mu$m with $N=100$). Here, we show that, for $10$ times smaller lattice spacing $a=10$ nm ($L=10$ $\mu$m with $N=1000$, $t\approx 27.21$ meV, $t_{\Gamma}=0.2t\approx 5.45$ meV), the key results are almost identical. As an example, we consider the setup with left lead being spin-up polarized and right lead being unpolarized and compare directly the corresponding nonlocal conductance map $G_{L(+)R_{(}N)}(\Delta_{Z},E)$ for the parameters $a=100$ nm ($N=100$) and $a=10$ nm ($N=1000$), see Fig. 14. Furthermore, we investigate whether the presence of disorder affects the results similarly when using two different lattice constants: $a=100$ nm ($N=100$) and $a=10$ nm ($N=1000$) while keeping the nanowire length constant at $L=10$ $\mu$m. In the latter case we set even stronger disorder $|\delta\mu_{i}|\lesssim 1$ meV (Gaussian distribution with $\sigma=0.3$ meV) to have similar mean free paths in both cases (see appropriate comment at the end of this Appendix). It can be observed in Fig. 14 (b,b’) that the nonlocal conductance signature of the Andreev band spin remains clearly visible also for the smaller lattice constant. For completeness, we also calculate maps of local $G_{LL}(\Delta_{Z},E)$ and nonlocal conductance $G_{LR}(\Delta_{Z},E)$ for other configurations of spin polarized leads: L(+)R(-), L(-)R(+), L(+)R(+), L(-)L(-) which are illustrated in Fig. 15 for lattice spacing $a=10$ nm ($L=10$ $\mu$m with $N=1000$, $t\approx 27.21$ meV, $t_{\Gamma}=0.2t\approx 5.45$ meV) and disorder amplitude $|\delta\mu_{i}|\lesssim 1$ meV. In order to estimate the mean free path $l_{mfp}$ induced by the onsite electrostatic disorder $\delta\mu_{i}$ one can use Fermi’s golden rule which gives Thamm_2024 ; Cole_2016 ; Awoga_2023 :
$l_{mfp}=v_{F}\tau\simeq\frac{3\hbar v_{F}}{2\pi aN_{1D}(0)W^{2}}$ (12)
where $N_{1D}(0)=2/(\pi\hbar v_{F})$ is the 1D density of states, and W the disorder strength corresponding to the amplitude of fluctuations of the onsite chemical potential $-W<\delta\mu_{i}<W$. This yields the following estimates for the mean free path: $l_{mfp}\simeq 685$ nm ($l_{mfp}\simeq 617$ nm) for $a=100$ nm and $W=0.3$ meV ($a=10$ nm and $W=1$ meV).
Figure 15: The nonlocal conductance maps for various configurations of spin polarization in the leads are shown in the following panels: (a) $G_{L(+)R(-)}(\Delta_{Z},E)$, (a’) $G_{L(-)R(+)}(\Delta_{Z},E)$, (a”) $G_{L(+)R(+)}(\Delta_{Z},E)$, (a”’) $G_{L(-)R(-)}(\Delta_{Z},E)$. The corresponding local conductance maps $G_{L(\pm)L(\pm)}(\Delta_{Z},E)$ are shown on the lower panels (b-b”’). The onsite disorder is presented in panel (c), while the spin and charge densities for selected states, marked by blue and red symbols, are depicted in panels (d) and (e), respectively, similar to Fig. 11 in the main text.
Figure 16: The $x$ component of spin the density $S_{x}(x)$ as a function of position along the nanowire for the lowest and second positive energy states are shown in panels (a) and (b), respectively. The primed panels (a’), (b’) show a zoom-in of the $S_{x}$ oscillations at the left end of the nanowire. Colors denote different values of the spin-orbit strength $\alpha$ (as indicated by the legend), which affects the amplitude and frequency of oscillations, as well as the position where the spin density changes sign from positive to negative. Here, we set $L=9$ $\mu$m, $a=30$ nm, $N=300$, $\Delta_{Z}=0.4$ meV, $\Delta_{sc}=0.35$ meV.
Appendix D Spin density profile at nanowire ends
Figure 17: Same as Fig. 16 but for larger Zeeman energy set to $\Delta_{Z}=0.45$ meV which results in a larger induced gap ($\Delta_{i}=2|\Delta_{sc}-\Delta_{Z}|=0.2$ meV). This larger induced gap suppresses the amplitude of spin density $S_{x}(x)$ oscillations at the nanowire ends.
Here, we examine how the system parameters - Zeeman energy $\Delta_{Z}$, spin-orbit strength $\alpha$, and length of the system $L$ affect the spin density profile $S_{x}(x)$ at the ends of the nanowire for the two lowest positive energy bulk states. To be specific, we first set the parameters to $L=9$ $\mu$m, $a=30$ nm, $N=300$, $\Delta_{sc}=0.35$ meV, $\Delta_{Z}=0.4$ meV such that the system is near the TPT and the induced gap is small. We then plot the spin densities $S_{x}(x/a_{x})$ along the wire for the first two energetically lowest states (with positive energy) varying the spin-orbit strength from $\alpha=35$ meVnm to $\alpha=60$ meVnm (see Fig. 16). One can observe that both the amplitude and frequency of the spin density oscillations increase with $\alpha$. Next, we increase the Zeeman energy to $\Delta_{Z}=0.45$ meV which moves the system further away from the TPT and makes the induced gap larger. As shown in Fig. 17 the amplitude of $S_{x}(x)$ oscillations is smaller than in the previous case. In the subsequent analysis, we fix the spin-orbit strength to $\alpha=40$ meVnm and vary the Zeeman energy between $\Delta_{Z}=0.38$ meV and $\Delta_{Z}=0.52$. The corresponding Fig. 18 demonstrates that the amplitude, frequency, and the position where the spin density changes sign are strongly dependent on $\Delta_{Z}$. We expect that the oscillation and sign change of the spin density, which occur only in the topological phase, reflect the presence of a zero-energy Majorana state in the system’s energy spectrum, whose wave function oscillates around zero at the nanowire end. Most likely, these oscillations arise from the interplay between interior and exterior energy gaps composite_majorana . Finally, we investigate how the length of the system affects spin density oscillations. Fig. 19 shows that the length of the system $L$ primarily influences the amplitude of the oscillations: the amplitude increases as the length decreases. This effect is likely related to the hybridization of Majorana states, which becomes more pronounced with a shorter wire. The observation that the spin density changes sign near the end only in the topological phase is probably due to the presence of Majorana states in the energy spectrum, which also affects the spin density of other energy states. This behavior could potentially serve as a new hallmark of the topological superconducting phase, however, a more detailed analysis is beyond the scope of this work and could be a topic for future research.
Figure 18: Similar to Figs. 16 and 17 but for different values of Zeeman energy ranging from $\Delta_{Z}=0.38$ meV to $\Delta_{Z}=0.52$ meV. The spin-orbit strength is fixed at $\alpha=40$ meVnm. The colors corresponds to the spin densities $S_{x}(x)$ plotted for different values of Zeeman energy $\Delta_{Z}$ (as indicated by legend).
Figure 19: Similar to previous figures Figs. (17-18) but for different lengths of the system ranging between $L=4.5$ $\mu$m to $L=12$ $\mu$m. The spin-orbit strength is fixed at $\alpha=40$ meVnm and the Zeeman energy to $\Delta_{Z}=0.45$ meV. The colors corresponds to the spin densities $S_{x}(x)$ for these different system lengths. One can clearly see that the amplitude of spin density oscillations decreases as the length of the system increases.